OpenStudy (anonymous):

The sum of four numbers in A.P is 20. The ratio of the product of the first and the last numbers and the product of 2nd & 3rd numbers is 2:3 ; Find the numbers . . . ANSWER = 2,4,6,8 OR 8,6,4,2

OpenStudy (anonymous):

Cant understand how to start

OpenStudy (anonymous):

Let four numbers in AP be a-3d ,a-d ,a+d and a+3d

OpenStudy (anonymous):

Now, a-3d+a-d+a+d+a+3d=20 4a=20 a=5

OpenStudy (anonymous):

Also, (a-3d)(a+3d)/ {(a-d) (a+d)} =2/3

OpenStudy (anonymous):

i cant understand this sentence "The ratio of the product of the first and the last numbers and the product of 2nd & 3rd numbers is 2:3"

OpenStudy (anonymous):

It simply means (a-3d)(a+3d)/ {(a-d) (a+d)} =2/3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

what should i do next.. cross multiply?

OpenStudy (anonymous):

Now, 3(a-3d)(a+3d) =2(a-d)(a+d) 3a^2 -27d^2 =2a^2 -2d^2 a^2 =25d^2 Also a=5 5^2 =25d^2 d^2=1 d=+-1

OpenStudy (anonymous):

Now, can u?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have to substitute the values of a and d in the four numbers which were a-3d, and so on

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok got the numbers :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome